40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Oct 18 17:35:07 2019, gautam, Update, SUS, ETMX sat box disconnected 
    Reply  Mon Oct 21 16:02:21 2019, gautam, Update, SUS, ETMX over the weekend filteredData.pdf
Message ID: 14982     Entry time: Mon Oct 21 16:02:21 2019     In reply to: 14977
Author: gautam 
Type: Update 
Category: SUS 
Subject: ETMX over the weekend 

Looking at the sensor and oplev trends over the weekend, there was only one event where the optic seems to have been macroscopically misaligned, at ~11:05:00 UTC on Oct 19 (early Saturday morning PDT). I attach a plot of the 2kHz time series data that has the mean value subtracted and a 0.6-1.2 Hz notch filter applied to remove the pendulum motion for better visualization. The y-axis calibration for the top plot assumes 1 ct ~= 1 um. This "glitch" seems to have a timescale of a few seconds, which is consistent with what we see on the CCD monitors when the cavity is locked - the alignment drifts away over a few seconds.

As usual, this tells us nothing conclusive. Anyways, I am re-enabling the watchdog and pushing on with locking activity and hope the suspension cooperates.

Quote:
 

The satellite box was reconnected and the suspension was left with watchdog off but OSEM roughly centered. We will watch for glitches over the weekend.

Attachment 1: filteredData.pdf  187 kB  Uploaded Mon Oct 21 17:05:17 2019  | Hide | Hide all
filteredData.pdf
ELOG V3.1.3-